OpenStudy (anonymous):

if a wire of resistance 36 ohm is bend in the form of circle how will we find effective resistance between A and B which make sector of angle 30

OpenStudy (anonymous):



OpenStudy (anonymous):

Answer is 4.83 R1 - R2 = 36-36Cos30=Answer

OpenStudy (ybarrap):

The fraction of the wire between A and B is $$ {30\times\cfrac{\pi}{180}\over 2\pi }=\cfrac{1}{12} $$ So the fraction of resistance in that part of the wire is $$ \cfrac{1}{12}\times 36=3 ~\Omega $$

OpenStudy (radar):

However the resistance will consist of 3 Ohm in parallel with 33 OhmSo the effective resistance is 2.75 Ohm